A survey showed that 73​% of adults need correction​ (eyeglasses, contacts,​ surgery, etc.) for their eyesight. If 22 adults are randomly​ selected, find the probability that at least 21 of them need correction for their eyesight. Is 21 a significantly high number of adults requiring eyesight​ correction? Please show ti83 steps
Answer #1

Ans:

Use binomial distribution with n=22 and p=0.72

P(Atleast 21)=P(x>=21)

=P(x=21)+P(x=22)

=22C21*0.7321*(1-0.73)1+0.7322

=0.0080+0.0010

=0.0090

As,above probability is less than 0.05,so,yes,21 a significantly high number of adults requiring eyesight​ correction.

*For ti83,Use P(x>=21)=1-P(x<=20)=1-binomcdf(22,0.73,20)=0.0090